OpenStudy (anonymous):

If the width of a rectangle is 16 inches less than twice the length and the area is 96 square inches, find the width of the rectangle.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

let the length be x. so width= 2x-16 area = length * width= (x)(2x-16)=96 2x^2-16x-96=0 x^2-8x-48=0 x=12. length =12, width= 8

OpenStudy (anonymous):

thanks